# Black Holes as Condensation Points of Fuzzy Dark Matter Cores

### v1

@@ Abstract
Summary of results. From our simulations we find that prior to the core condensation, miniclusters of FDM can form, that would collapse in themselves and condensate as shown in Levkov et al. (2018); Chen et al. (2021) and in Appendix A. However the presence of the BH modifies such evolution and the mini-cluster merges with the black hole leading to the following effects. First, the FDM mass accommodates around the black hole, which shows how the black hole acts as an attraction point even for the smallest black hole mass case we used.
@@
Second, the motion of the BH is oscillatory, a behavior found in Boey et al. (2024) while studying the dynamical friction of FDM on black holes using formed spherically symmetric cores; here we showed that the oscillatory behavior happens even prior to core-condensation, within a non-spherically symmetric, time-dependent granular FDM distribution. The reheating of the BH motion due to the dynamics of the FDM can be observed as well.
@@
Third, the analysis of $\rho_{max}$ as function of time reveals that the FDM cloud starts condensing once it is around the BH, however its growth is affected by the presence of the BH motion. The motion of the BH within the core disperses the central density away, reducing the central averaged density of the FDM core. We find smaller central density of FDM for a bigger BH mass that drags the FDM around with its oscillatory motion.
@@
Fourth, the solid angle and time averaged density of FDM around the black hole, can be modeled with the stationary spherically symmetric solutions of the FDM+BH eigenvalue problem. The implication is that stationary solutions act like attractor solutions of FDM in the presence of a Black Hole.
In summary, our analysis shows that Black Holes act as condensation points where kinetic relaxation can take place. We consider that the influence of the black hole in the galaxy formation context can lead to consistency checks of the FDM model, particularly in scenarios of galactic cores with supermassive black holes at their centers. The effects described in this work can lead to new detectable predictions that in turn should challenge the viability of the FDM model.
@@ Acknowledgments
We thank Pierre-Henri Chavanis for providing useful comments. Curicaveri Palomares and Iván Álvarez receive support from the CONAHCyT graduate scholarship program. This research is supported by grants CIC-UMSNH-4.9, Laboratorio Nacional de Cómputo de Alto Desempeño Grant No. 1-2024, CONAHCyT Ciencia de Frontera 2019 Grant No. Sinergias/304001.
@@ Appendix A Pure FDM
As a test we show the core condensation without a BH in Figure 6, that contains snapshots of the density at different times on a plane that passes through the point of maximum density, where we can notice that between $t\approx 10$ and $t\approx 30$ a minicluster is formed and in later times the solitonic core gets condensed. Further analysis that evidences condensation appears in Figure 7 where the maximum density is plotted as function of time in units of the condensation time $\tau_{g}\propto\frac{v^{6}}{\bar{\rho}^{2}\log{(vL)}}$ as found in Chen et al. (2021), which for our simulation corresponds to $\tau_{g}\approx 7$. The results obtained are in agreement with the findings in Chen et al. (2021) related to density growing with a power law of $t$.
@@
Figure 7: Maximum density $\rho_{max}$ in the numerical domain as function of time in units of $\tau_{g}$. This plot illustrates the condensation process as shown in Levkov et al. (2018), which establishes the time scale $\tau_{g}$ for condensation, indicated with the vertical line.
Appendix B Stationary Solutions
B.1 Solution of the FDM+BH eigenvalue problem
@@
where $V_{\bullet}=-GM_{BH}/r$ is the gravitational potential due to a black hole of mass $M_{BH}$. The problem is solved in code units, spherical symmetry and harmonic time dependence of the parameter order $\Psi=\psi(r)e^{-iwt}$ with $\psi$ a real function. These assumptions lead to an eigenvalue problem for the eigenvalue $\omega$, provided isolation boundary conditions at infinity and regularity at the origin.
We note that this system exhibits invariance under the transformation $\{t^{\prime},\Psi^{\prime},x^{\prime},V^{\prime},\rho^{\prime},M_{BH}^{\prime}\}~{}\rightarrow\{\lambda^{-2}t,\lambda^{2}\Psi,\lambda^{-1}x,\lambda^{2}V,\lambda^{4}\rho,\lambda M_{BH}\}$,where $\lambda$ is an arbitrary parameter. Now, for the construction of a phenomenological model of the FDM density we define the invariant $\alpha=M_{BH}^{2}/\psi_{0}$ under the $\lambda-$transformation, that allows one to parametrize the family of solutions of the eigenvalue problem. We note that in the limit $V_{\bullet}\gg V$ or equivalently $\alpha\gg 1$, the system reduces to the case of the hydrogen atom with ground state solution $\psi(r)=2\left(\frac{1}{a_{0}}\right)^{3/2}e^{-r/a_{0}}$, being $a_{0}$ the Bohr radius. We thus propose a formula that in the limit of high $\alpha$ approaches the exponential solution
@@
where $\rho_{c}$ is the central density, the core radius $r_{c}$ is defined as the radius where the density decreases to the half of its central value and $\beta$ is an $\alpha$ dependent function to be found.

### v3

@@ IV Conclusions
From our simulations we find that prior to the core condensation, miniclusters of FDM can form, that would collapse in themselves and condensate as shown in Levkov et al. (2018); Chen et al. (2021) and in Appendix A. However the presence of the BH modifies such evolution and the mini-cluster merges with the black hole leading to the following effects. First, a rather expected result between two attracting objects, the FDM gas density distributes around the black hole, despite the precollapsed minicluster is beyond the BH’s tidal radius or not, which shows how the black hole acts as an attraction point even for the smallest black hole mass case used.
@@
Second, the motion of the BH with respect to the center of the core is oscillatory, a behavior found in Boey et al. (2024) while studying the dynamical friction of FDM on black holes using formed spherically symmetric cores; here we showed that the oscillatory motion happens even prior to core-condensation, within a non-spherically symmetric, time-dependent granular FDM distribution. The reheating of the BH motion due to the dynamics of the FDM can be observed as well.
@@
Third, the analysis of $\rho_{max}$ as function of time reveals that the FDM cloud starts condensing once it is around the BH, however its growth is affected by the presence of the BH motion. The motion of the BH within the core disperses the central density away, reducing the averaged central density of the FDM core. We find smaller central density of FDM for a bigger BH mass, which can be understood in terms of the relative motion of the BH with respect to the core. This effect is contrasted with the test-field case, in which the BH does not back-react to the gravity of the FDM and does not scatter the FDM around.
@@
Fourth, we have developed a spherically symmetric phenomenological formula for the FDM density that can be used to fit FDM cores around BHs. The accuracy of this model suggests that stationary solutions act like attractor solutions of FDM in the presence of a Black Hole.
In summary, our analysis shows that Black Holes act as condensation points where kinetic relaxation can take place. We consider that the influence of the black hole in the galaxy formation context can lead to consistency checks of the FDM model, particularly in scenarios of galactic cores with supermassive black holes at their centers. The effects described in this work can lead to new detectable predictions that in turn should challenge the viability of the FDM model. Potential scenarios to be analyzed include the motion of the BH relative to the core, for example, the oscillations seen in Fig. 3 for the smallest black hole have a period of order 14Myr, that eventually could be correlated to the variability near SMBHs, and its frequency should be correlated with the granularity of the FDM distribution for a statistically significative number of initial conditions. In this same context heating and reheating when the BH mass is small, near the test particle regime, can lead to chaotic motion within FDM cores Alvarez-Rios and Guzman (2025) and in turn contribute with wandering BHs within cores. Another potential application is related to the evidence that the presence of the BH affects the averaged central FDM density as seen in Fig. 2, which allows the comparison with the systematic modeling of dark matter distribution at M87 De Laurentis and Salucci (2022), using also luminous matter models coupled to FDM Alvarez-Rios et al. (2024).
@@ Acknowledgments
We thank Pierre-Henri Chavanis for providing useful comments. Curicaveri Palomares and Iván Álvarez receive support from the CONAHCyT graduate scholarship program. This research is supported by grants CIC-UMSNH-4.9, Laboratorio Nacional de Cómputo de Alto Desempeño Grant Nos. 1-2024 and 5-2025, CONAHCyT Ciencia de Frontera 2019 Grant No. Sinergias/304001.
@@ Appendix A Pure FDM
As a test we show the core condensation without a BH in Figure 6, that contains snapshots of the density at different times on a plane that passes through the point of maximum density, where we can notice that between $t\approx 10$ and $t\approx 30$ a minicluster is formed and in later times the solitonic core gets condensed.
The initial conditions in Fourier space and integrated mass, correspond to a standardized FDM set up in Chen et al. (2021); we use this Appendix to verify that our code reproduces the results of such standard simulation. Due to the randomness of the initial conditions, the collapse of the minicluster s the result of a random fluctuation in a random location of the domain, where the configuration later collapses. The condensation process appears in Figure 7 where the maximum density is plotted as function of time in units of the condensation time $\tau_{g}\propto\frac{v^{6}}{\bar{\rho}^{2}\log{(vL)}}$ as found in Chen et al. (2021), which for our simulation corresponds to $\tau_{g}\approx 7$. The results obtained are in agreement with the findings in Chen et al. (2021) related to density growing with a power law of $t$.
@@
Figure 7: Maximum density $\rho_{max}$ in the numerical domain as function of time. This plot illustrates the condensation process as shown in Levkov et al. (2018), which establishes the time scale $\tau_{g}$ at $t\sim 7$ for condensation, indicated with the vertical line.
What is shown in our analysis is that when a BH is introduced within the FDM initial conditions, the condensation process develops at the BH, even if the minicluster formed in a random location has to be first disrupted and the redistributed around the BH.
Appendix B Stationary Solutions of the FDM+BH eigenvalue problem
@@
where $V_{\bullet}=-GM_{BH}/r$ is the gravitational potential due to a black hole of mass $M_{BH}$. The problem is solved in code units, spherical symmetry and harmonic time dependence of the parameter order $\Psi=\psi(r)e^{-iwt}$ with $\psi$ a real function. These assumptions lead to the following eigenvalue problem for the eigenvalue $\omega$, provided isolation boundary conditions at infinity and regularity at the origin:
$w\psi$ $=$ $\left[-\dfrac{1}{2r^{2}}\frac{d}{dr}\left(r^{2}\frac{d}{dr}\right)+\left(V+V_{\bullet}\right)\right]\psi,$ (9)
$\frac{1}{r^{2}}\frac{d}{dr}\left(r^{2}\frac{d}{dr}\right)V=\left(\rho-\bar{\rho}\right).$
We rewrite this problem with a set of two first order equations by defining the variables $\phi=r^{2}\psi^{\prime}$ and $M=r^{2}V^{\prime}$, where denotes derivative along the radial direction. The eigenvalue problem as aa first order system reads:
$\frac{d\psi}{dr}$ $=$ $2r^{2}(-w+V+V_{\bullet})\psi,$
$\frac{dM}{dr}$ $=$ $r^{2}(\rho-\bar{\rho}),$
$\frac{d\psi}{dr}\psi$ $=$ $\frac{\phi}{r^{2}},$
$\frac{dV}{dr}$ $=$ $\frac{M}{r^{2}},$ (10)
that we solve using the shooting method with the boundary conditions at the origin $\psi(0)=0,M(0)=0,\psi(0)=\psi_{0},V(0)=V_{0}$, and at an external boundary $\psi(r\rightarrow\infty)=0,\phi(r\rightarrow\infty)=0$ and $V(r\rightarrow\infty)=0$, being $\psi_{0}$ a fixed central value; then the potential is rescaled to impose a monopolar boundary condition at the external boundary.
We note that this system exhibits invariance under the transformation $\{t^{\prime},\Psi^{\prime},r^{\prime},V^{\prime},\rho^{\prime},M_{BH}^{\prime}\}~{}\rightarrow\{\lambda^{-2}t,\lambda^{2}\Psi,\lambda^{-1}r,\lambda^{2}V,\lambda^{4}\rho,\lambda M_{BH}\}$,where $\lambda$ is an arbitrary parameter. Now, for the construction of a phenomenological model of the FDM density we define the invariant $\alpha=M_{BH}^{2}/\psi_{0}$ under the $\lambda-$transformation, that allows one to parametrize the family of solutions of the eigenvalue problem. In order to cover a wide parameter space, we solve the system for values of $\alpha$ within the range $\alpha\in[0,500]$, and the results are shown in Figure 8, where the $\alpha=0$ is the ground state soliton solution for the SP system corresponding to the case without black hole.
Figure 8: Solution densities obtained from the numerical solution of the eigen-value problem (10) for $\alpha$ in the range $[0,500]$. The figure shows that the FDM density becomes more compact as $\alpha$ increases.
We note that in the limit $V_{\bullet}\gg V$ or equivalently $\alpha\gg 1$, the system reduces to the case of the hydrogen atom with ground state solution $\psi(r)=2\left(\frac{1}{a_{0}}\right)^{3/2}e^{-r/a_{0}}$, being $a_{0}$ the Bohr radius. This exponential behavior can be seen in the numerical solution for high $\alpha$ in Figure 8, thus we propose a formula that in the limit of high $\alpha$ approaches the exponential solution
@@
where $\rho_{c}$ is a central density, a core radius $r_{c}$ is defined as the radius where the density decreases to the half of its central value and $\beta$ is an $\alpha$, dependent function to be found.
@@
Evolution of stationary solution. Now that we have a model for the density profile of a stationary solution we test the stability of configurations constructed in this way. Since the formula does not exactly give the profile obtained from the solution of the eigenvalue problem, the one constructed with the formula can be considered as a solution to the eigenvalue problem plus a perturbation. Then, the effects of the perturbation will trigger oscillation modes of the equilibrium configuration.
We now show that these solutions of the SP system with a black hole in the center are stable, and that the code works in this scenario. For this we evolve some configurations with different values of $\alpha$. The evolution is carried out in the numerical cubic domain of side $L=20$ with the black hole at the center, using resolution $\Delta x=20/128$. The model for the black hole potential uses the prescription in Eq. (5).
The results of the evolution of various configurations are shown in Figure 10 and can be compared with the numerically perturbed solitons in Guzmán and Ureña López (2006). Likewise in the pure soliton solution Guzmán (2019), under a perturbation the core oscillates with a dominant frequency mode.
Figure 10: Oscillation of the central density for configurations with different $\alpha$.
Appendix C Test-field condensation of FDM
Prior to the main scenario involving an interacting BH with a fluctuating medium of FDM, there is an intermediate scenario in which the BH remains fixed and does not respond to the fluctuating gravitational field sourced by the FDM. This allows us to track the condensation process due to the sole presence of the black hole, without reheating its motion and without the BH scattering the FDM around.
We use the same initial conditions and black hole masses as in the fully coupled case. The evolution is illustrated in Figure 11. The condensation process is monitored with the maximum density as function of time in Figure 12. This result is to be contrasted with that in Figure 2.
Figure 11: Snapshots of the density at different times for $M_{BH}=M/256,~{}M/128,~{}M/64,~{}M/32$, before and after the black hole captures the minicluster.
Figure 12: Evolution of $\rho_{max}$ in simulations with $M_{BH}=M/256,M/128,M/64,M/32$ in the test field regime. Unlike in the case where the BH reacts to the FDM gravitational potential, the maximum density does not decrease.